OpenStudy (anonymous):

Write the equation of a line, in general form, with a point (-3, 0) and an undefined slope?

OpenStudy (anonymous):

use Y-Y1=m(X-X1) leave m as it is and X1,Y1 is a point in on the line that is (-3,0)

OpenStudy (anonymous):

y-3=0?

OpenStudy (anonymous):

no y-0=m(x-(-3)) simplify it you will get y=mx+3m which is the answer

OpenStudy (anonymous):

oh ok thanks